i'm a freshman i college not real good with math..just wanted to know if i was on the right track..9v=15...i came up with 6/1..

Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

5√75 – √48 + 8√3
= 5sqrt(25*3) - sqrt(16*3) + 8sqrt3
= 5*5sqrt3 -4sqrt3 + 8sqrt3
= (25 - 4 + 8)sqrt3
=29sqrt(3)
You are correct.
29√3 I think is the answer, if so can it be simlified even more?
Cannot be further simplified.
=====================================
59 – 33 ÷ (13 - 4²) – 8
= 59 - 33 /(-3) -8
= 59 + 11 -8
=62
================
multiplication and division must be done before addition and subtraction.
================
Cheers,
Stan H.

Answer by scott8148(6628)

first one looks good

remember PEMDAS __ 59–33÷(13-4²)–8 __ 59–33÷(-3)–8 __ 59-(-11)-8 __ 70-8 __ 62